OpenStudy (anonymous):

Make up an equation of the form ax+b=c that has -3 as it's solution

OpenStudy (anonymous):

If c=-3, a=2, x=2 and b=-7 Therefore, 2(2)+(-7)=-3.

OpenStudy (anonymous):

Awesome thanks

OpenStudy (cwrw238):

if you mean an equation whose solution is x = -3 that is not the answer

OpenStudy (anonymous):

The question was: stated correctly that has -3 as it solution.

OpenStudy (cwrw238):

the equation that jaeriko wrote has 2 as its solution

OpenStudy (anonymous):

The problem is asking to write a equation of the form ax+b=c with the solution being -3. I thought his answer works out as 2x2= 4+ -7= -3. I am new at algebra but when so I could be wrong so is this not correct?

OpenStudy (cwrw238):

the question is a bit ambiguous usually when you talk abour the solution of an equation you mean the value of x which fits the equation for example solve 2x + 3 = 15 solving 2x + 3 - 3 = 15 - 3 2x = 12 2x/2 = 12/2 x = 6 6 is the solution

OpenStudy (anonymous):

That helps me but the question is stated in the assignment as is so I'm not sure what else to state thanks for your input

OpenStudy (cwrw238):

ok if you want an answer in line with my above input it would be 2(-3) + 10 = 4 so the equation is 2x + 10 = 4

OpenStudy (anonymous):

Thank you for your help

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Oh, was x supposed to equal -3? I thought he meant an equation that would equal -3, not one that used -3 as its x value.

OpenStudy (cwrw238):

yes that s my point - a solution of an equation is the value of x not the value of the function of x

OpenStudy (anonymous):

I'm new at this but the way it reads is -3 is the solution of noted equation.

OpenStudy (cwrw238):

then it is referring to the solution of the equation so 2x + 10 = 4 would be one such equation ax + b = c is referred to as a general form of an equation x is the variable to be found and a,b and c refer to constant values

OpenStudy (anonymous):

So what does the -3 have to do with the solution. Your explanation helps, it's one that was brought up in class

OpenStudy (cwrw238):

well x = -3 is the solution of the equation 2x + 10 = 4 which means that if we insert x = -3 into the equation then it balances that is left hand side = right hand side its said that x=-3 'satisfies' the eqaution.

OpenStudy (anonymous):

Ohh got it you have been very helpful